OpenStudy (balala):

Could somebody please help me with this project? I tried to do it but I'm not sure I did it right? Could you just check it or something please?

OpenStudy (balala):

My project is attached here:

OpenStudy (balala):

@arabpride could you help? :)

OpenStudy (arabpride):

I'm not even sure... did you plug those numbers in those graphs?

OpenStudy (balala):

Um I did all the work that is colored blue :/

OpenStudy (arabpride):

Well I'm positive the input and the output is correct, however, I've never done something like this, so I'm not sure whether you've completed it ENTIRLEY or not... think so tho /.^

OpenStudy (arabpride):

are correct*

OpenStudy (balala):

alrighty... thanks so much for taking the time to check!

OpenStudy (arabpride):

No Problem!! *^.^* ~Good Luck!!!!

OpenStudy (aum):

The first column gives ingredient for FOUR SERVINGS. If we assume each serving is for ONE PERSON, then we will need to find the amount of ingredient needed for 24 servings. So the first row might look like this: 4 skinless, boneless 4:4 = 24:x x = 24 * 4 / 4 = 24 chicken breasts .... ¼ cup all-purpose 4:¼ = 24:x x = 24 * ¼ / 4 = 24 / 16 = 1.5 flour ....